OpenStudy (anonymous):

How many electrons are in 2.86kg of water?

OpenStudy (turingtest):

this is chemistry, but here we go...

OpenStudy (jamesj):

Nice question

OpenStudy (turingtest):

I think I remember how to do these, tell me if I mess up...

OpenStudy (jamesj):

(find the molar mass of water; find how many moles of water 2.86 kg is; multiply by Av. number to find the # of molecules; multiply that by the # electrons per molecule)

OpenStudy (turingtest):

\[(2.86\times10^3g)(\frac{mol}{18.02g})(\frac{6.02\times 10^{23}molecules}{mol})(\frac{10e}{molecule})\]right?

OpenStudy (jamesj):

looks right

OpenStudy (turingtest):

Chem was the only AP I got a 3 on :( but stoichiometry is pretty easy. I got lost by titrations though

OpenStudy (anonymous):

Thank you, I know it appears to be chemistry but it is part of my physics 202 HW for Electricity.

OpenStudy (turingtest):

Fair enough, makes sense it could be in physics I suppose.

OpenStudy (jamesj):

I'm happy to have it the question here! ;-) Everyone who studies physics should understand some basic chemistry anyway.